# Selecting informative conformal prediction sets with false coverage rate control

### v1

@@ E.1 Excluding a null value $y_{0}\in\mathbb{R}$
An illustration is provided in Figure 10 in the non-parametric Gaussian model for different situations: Rows 1,2 correspond to an homoscedastic model with perfect prediction of the variance and a mean predictor with various accuracy (less accurate in the middle for row 1, more accuracy in the middle for row 2). Rows 3,4 correspond to an heteroscedastic model with perfect prediction of the mean and a the variance predictor under-estimating the variance in the middle for row 3, and over-estimating the variance in the middle for row 4). The marginal prediction interval (no selection) is displayed in light blue while the prediction interval after selection is displayed in dark red. The quantities reported at the top of each panel are the FCR and adjusted power averaged over $1000$ repetitions (while each panel only displays the last experiment as a typical realisation of the sample).
@@
Figure 10: Informative selection by excluding the null value $y_{0}$ in the (homoscedastic or heteroscedastic) regression case, see § E.1.
First, as expected, the naive classical conformal selection does not provide a correct FCR control in all cases, while InfoSP and InfoSCOP do control the FCR in any case. Second, InfoSCOP always improves InfoSP in terms of power, and the range of improvement depends on where non-covering errors are likely to happen: when errors are more likely to arise on the selection (rows 1,3), the behavior of InfoSP and InfoSCOP are similar; when errors are less likely to arise on the selection (rows 2,4), InfoSCOP improves over InfoSP in a striking manner (and is even better than classical conformal). This is both because of the reduction of the selection effect and because the pre-processed $p$-values are calibrated with much lower residuals and thus are much more efficient than the original $p$-values.
E.2 Prediction intervals for Jin and Candes, (2023) selection
@@
We consider here the aim of finding one-sided prediction intervals on the selected. Let us assume that the score function is monotone in the following sense (Jin and Candes,, 2023): for all $x\in\mathbb{R}^{d}$, for $y\leq y^{\prime}$, $S_{y}(x)\geq S_{y^{\prime}}(x)$. Examples of monotonic scores are given by $S_{y}(x)=(\mu(x)-y)/\sigma(x)$, $S_{y}(x)=(\mu(x)-My)/\sigma(x)$ with $M=2\sup_{x}\mu(x)$. The following result summarizes our finding in this case.
Corollary E.1.
Consider the iid model in the regression case and assume that the score function is monotone (see above) and such that Assumption 2 (ii) (iii) and Assumption 1 hold. Then the following holds for InfoSP with informative collection =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>I\cap(-\infty,y_{0}]=\emptyset\}$ and $p$-value collection $\mathbf{\overline{p}}$ (7):
@@
    $q_{i}=\overline{p}^{(y_{0})}_{i}=\frac{1}{n+1}\Big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})\geq S_{y_{0}}(X_{n+i})\}}\Big{)}.$
@@
Importantly, (iii) ensures that the produced prediction intervals are always “doing what they should do”, by always excluding values smaller than $y_{0}$.
@@
Obviously, a similar result holds for InfoSCOP, for any initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ that satisfies the permutation preserving Assumption 3. Corollary E.1 is illustrated on Figure 11 when $\mathcal{S}^{(0)}$ is taken here has $\mathrm{BH}(\mathbf{q})$ at level $2\alpha$ with the score $S_{y}(x)=(\mu(x)-y)/\sigma(x)$. The comments are qualitatively similar to those of Figure 10: when covering errors are less likely on the selection, the improvement of InfoSCOP over InfoSP is substantial.
@@
Figure 11: Same as Figure 10 with Jin and Candes, (2023) type selection and one-sided prediction intervals, see § E.2.
E.3 Comparison with existing selective prediction sets
@@
Figure 12: Comparison to previous conformal selective inferences in the regression case: new informative prediction intervals (InfoSCOP) versus non-informative prediction intervals (SCOPab) (left: homoscedastic Gaussian regression with perfect variance prediction and errors in the mean prediction; right: heteroscedastic Gaussian regression with perfect mean prediction and errors in the variance prediction). Informative means here prediction intervals that does not contain $y_{0}$ (dashed line), see § E.3.
@@
Figure 12 is useful to visualize the difference between the selection proposed by Bao et al., (2024) and our informative selection. We display selective prediction intervals for two FCR controlling selections proposed in Bao et al., (2024): SCOPa uses a thresholding rule $\mathcal{S}=\{i\in[m]\>:\>\mu(X_{n+i})\geq y_{0}\}$, while SCOPb selects the largest $\mu(X_{n+i})$’s. Each time, a suitable conditional conformal prediction set is built in Bao et al., (2024) and reported in Figure 12 in green and purple. As one can see, either the selection is too conservative, or the prediction interval could include the nominal $b_{0}$. This not the case of InfoSCOP which always exclude $y_{0}$ by essence.
Appendix F Additional illustrations in the classification case
F.1 Three classes, each a mixture of bivariate normals
@@
In Figure 13 we show that the qualitative conclusions with regard to the respective procedures are as in § 5.3. Interestingly, in the hardest setting at the bottom row we see that the power of InfoSCOP is even larger than classic conformal. This is because the initial selection step not only prevents paying too much for selection, but it can also improve the calibration set for computing the conformal $p$-values after selection. In this hard setting, there is a large improvement of the tail probability that matters for inference. We demonstrate this in one realization of the data generation in Figure 14, where we can see that the 0.95 quantile of the original calibration set is much larger than the 0.95 quantile of examples that remain in the calibration set after the initial selection step. This is because the data generation is such that examples in the central cloud tend to have larger nonconformity scores $S_{Y_{i}}(X_{i})$ than the examples that are not in the central cloud, and most of the examples from the central cloud were eliminated from the calibration sample with the initial selection step. So the remaining nonconformity scores after the initial selection step tend to be smaller. Since the nonconformity scores of the examples from the test sample are unchanged, the improved calibration set after the initial selection step results in $p$-values that tend to be smaller when testing the null group, and in particular there are many more $p$-value that are below the 0.05 level.
@@
Figure 13: Selecting informative prediction sets in the iid setting with a large overlap between three classes. FCR (left column), resolution-adjusted power (middle column) and the expected fraction of covering prediction sets (right column) versus SNR for: balanced classes for minimally informative prediction sets (first row) and for prediction sets excluding a null class (second row); unbalanced classes for minimally informative prediction sets (third row) and for prediction sets excluding the largest class (fourth row). Each class is a mixture of a common component and a unique component, see § F.1 for details. Based on 2000 data generations, 1000 data points were used for training, and $n=m=500$.

### v3

@@ I.1 Excluding a null value $y_{0}\in\mathbb{R}$
An illustration is provided in Figure 7 in the non-parametric regression Gaussian model for different situations: Rows 1,2 correspond to an homoscedastic model with perfect prediction of the variance and a mean predictor with various accuracy (less accurate in the middle for row 1, more accuracy in the middle for row 2). Rows 3,4 correspond to an heteroscedastic model with perfect prediction of the mean and a the variance predictor under-estimating the variance in the middle for row 3, and over-estimating the variance in the middle for row 4). The marginal prediction interval (no selection) is displayed in light blue while the prediction interval after selection is displayed in dark red. The quantities reported at the top of each panel are the FCR and adjusted power averaged over $1000$ repetitions (while each panel only displays the last experiment as a typical realisation of the sample).
@@
Figure 7: Informative selection by excluding the null value $y_{0}$ in the (homoscedastic or heteroscedastic) regression case, see § I.1.
First, as expected, the naive classical conformal selection does not provide FCR control in all cases, while InfoSP and InfoSCOP do control the FCR in any case. Second, InfoSCOP always improves InfoSP in terms of power, and the range of improvement depends on where non-covering errors are likely to happen: when errors are more likely to arise on the selection (rows 1,3), the behavior of InfoSP and InfoSCOP are similar; when errors are less likely to arise on the selection (rows 2,4), InfoSCOP improves over InfoSP in a striking manner (and is even better than classical conformal). This is both because of the reduction of the selection effect and because the pre-processed $p$-values are calibrated with much lower residuals and thus are much more efficient than the original $p$-values.
I.2 Prediction intervals for Jin and Candes, (2023) selection
@@
We consider here the aim of finding one-sided prediction intervals on the selected. Let us assume that the score function is monotone in the following sense (Jin and Candes,, 2023): for all $x\in\mathbb{R}^{d}$, for $y\leq y^{\prime}$, $S_{y}(x)\geq S_{y^{\prime}}(x)$. A classical example of monotonic score function is given by $S_{y}(x)=(\mu(x)-y)/\sigma(x)$. The following result summarizes our finding in this case.
Corollary I.1.
Consider the iid model in the regression case and assume that the score function is monotone (see above) and such that Assumption 2 (ii) (iii) and Assumption 1 hold. Then the following holds for InfoSP with informative collection =$\mathcal{I}=\{I\mbox{ interval of }\mathbb{R}\>:\>I\cap(-\infty,y_{0}]=\emptyset\}$ and $p$-value collection $\mathbf{\overline{p}}$ (7):
@@
    $q_{i}=\overline{p}^{(y_{0})}_{i}=\frac{1}{n+1}\Big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})\geq S_{y_{0}}(X_{n+i})\}}\Big{)},$
    which coincides with the rejection set of the procedure proposed in Jin and Candes, (2023).
@@
In other words, the above result complements the multiple testing procedure of Jin and Candes, (2023), by providing in addition FCR controlling informative prediction sets on the selected outcomes.
@@
Obviously, a similar result holds for InfoSCOP, for any initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ that satisfies the permutation preserving Assumption 3. Corollary I.1 is illustrated on Figure 8 when $\mathcal{S}^{(0)}$ is taken here has $\mathrm{BH}(\mathbf{q})$ at level $2\alpha$ with the score $S_{y}(x)=(\mu(x)-y)/\sigma(x)$. The comments are qualitatively similar to those of Figure 7: when covering errors are less likely on the selection, the improvement of InfoSCOP over InfoSP is substantial.
@@
Figure 8: Same as Figure 7 with Jin and Candes, (2023) type selection and one-sided prediction intervals, see § I.2.
I.3 Predicting gene expression from promoter sequences
Vaishnav et al., (2022) fit a convolutional neural network model on tens of millions of random promoter sequences with the goal of predicting the expression level of a particular gene induced by a promoter sequence. They then used the model’s predictions to study the effects of promoters. They evaluated the performance on 61150 labeled sequences that were not used in training the model. The 61150 pairs of predicted and measured expressions will be used as a basis for a simulation to evaluate the performance of classic conformal, InfoSP, and InfoSCOP.
Suppose an analyst considers $m=50$ new random promoter sequences. From these, the analyst is interested in the prediction intervals for promoter sequences that are under-expressed. Importantly, the range of plausible values (i.e., a two-sided prediction interval) are of interest to the analyst, since it is not enough to establish with confidence that the expression is below a certain value, but also to know that the expression is above zero and evaluate the plausible extent to which it exceeds zero.
Suppose the analyst has access to the expression and predicted expression of $n=200$ random promoter sequences. These examples will serve as the calibration sample.
We consider the score function $S_{y}(x)=|\mu(x)-y|$, where $\mu(x)$ is the predicted expression (more sophisticated score functions, such as the locally weighted or quantile based functions discussed in § 4, cannot be used since we only have access to $\mu(x)$ from the prediction machine). We define an example as informative if its prediction interval is entirely below the threshold $a=9$. Therefore, the $p$-value for testing the (random) null hypothesis that the example is not informative is
$q_{i}=\max_{y\geq a}\overline{p}^{(y)}_{i}=\overline{p}^{(a)}_{i}\times{\mathbf{1}\{\mu(X_{n+i})<a\}}+{\mathbf{1}\{\mu(X_{n+i})\geq a\}},$
see Corollary 4.1 with $b=+\infty$.
In order to assess our methods using the 61150 labeled native yeast promoter sequences, we randomly sample labeled and unlabeled data sets as follows. For each of $B=10000$ trials, we randomly sample $n+m$ promoters, out of which $n$ points serve as the labeled calibration data set and $m$ points as the unlabeled test data set. The results are summarized in Table 1. Due to selection of informative examples, the classic conformal procedure that constructs each prediction interval at level $0.91-\alpha=$ 0.9 fails to control the FCR at the $\alpha=0.1$ nominal level. InfoSP provides the required control, but at the cost of some non-negligible power loss. InfoSCOP applies an initial selection step in which examples from the test sample and the second half the calibration sample are forwarded to InfoSP, only if their conformal $p$-value, computed using the first half of the calibration sample, for testing the (random) null hypothesis that the example is not informative, is at most $\alpha$. InfoSCOP controls the FCR at level $\alpha$ and has only slightly reduced power compared to classic conformal.
CC InfoSP InfoSCOP
FCR 0.150 0.030 0.081
Power 16.9 12.6 15.8
Res-adjusted Power 2.81 1.73 2.37
Table 1: The FCR, power, and resolution-adjusted power, for the three procedures considered for the yeast application (see text). The resolution adjusted power is the expected number of discoveries divided by the length of the prediction interval (since the length is fixed in this analysis). Based on $B=10000$ repetitions.
Figure 9 shows the prediction intervals for a single data generation. InfoSCOP selects almost the same number of informative examples as classic conformal, but all prediction intervals constructed with InfoSCOP cover the true value. Moreover, the lower bounds are all above zero. In practice, following such an analysis, the analyst can use the list of prediction intervals from InfoSCOP to the next stage: knowing with confidence the prediction intervals for informative promoters for this set can help the future design of $n+m$ examples with even lower expressions (see Vaishnav et al., (2022) for uses for the design of promoters for low gene expression). There will be a lab cost of culturing the yeast of $n$ examples to actually measure the expression level for $n$ promoters to serve as the calibration data, but $m$ can be quite large.
Figure 9: The observed versus predicted values for yeast expression, as well as the classic conformal prediction intervals upper and lower bounds (red lines), and the InfoSCOP prediction intervals (purple vertical lines), for a single data generation (see text in § I.3 for details). The false coverage proportion for classic conformal is 0.176 and 0 for InfoSCOP. Classic conformal reports 17 informative examples, but three examples are not covered by their prediction interval. InfoSCOP reports 16 informative examples (a subset of the 17 reported by classic conformal, including the three examples with non-covering classic conformal prediction intervals).
Remark I.1.
The selection in the above analysis fits within the Jin and Candes, (2023) selection framework outlined in the previous section, but with a different score function than the scores suggested Jin and Candes, (2023), since our score function is selected to provide two-sided prediction intervals.
I.4 Comparison with existing selective prediction sets
@@
Figure 10: Comparison to previous conformal selective inferences in the regression case: new informative prediction intervals (InfoSCOP) versus non-informative prediction intervals (SCOPab) (left: homoscedastic Gaussian regression with perfect variance prediction and errors in the mean prediction; right: heteroscedastic Gaussian regression with perfect mean prediction and errors in the variance prediction). Informative means here prediction intervals that does not contain $y_{0}$ (dashed line), see § I.4.
@@
Figure 10 is useful to visualize the difference between the selection proposed by Bao et al., (2024) and our informative selection. We display selective prediction intervals for two FCR controlling selections proposed in Bao et al., (2024): SCOPa uses a thresholding rule $\mathcal{S}=\{i\in[m]\>:\>\mu(X_{n+i})\geq y_{0}\}$, while SCOPb selects the largest $\mu(X_{n+i})$’s. Each time, a suitable conditional conformal prediction set is built in Bao et al., (2024) and reported in Figure 10 in green and purple. As one can see, either the selection is too conservative, or the prediction interval could include the nominal $y_{0}$. This is not the case of InfoSCOP which always exclude $y_{0}$ by essence.
Appendix J Additional illustrations in the classification case
J.1 Three classes, each bivariate normal
For the iid model, as described in § 5.2, we demonstrate the initial selection step in infoSCOP for excluding a null class in one realization of the data generation in Figure 11.
For non-trivial classification, as described in § 5.3, we provide numerical results in Figure 12.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
J.2 Three classes, each a mixture of bivariate normals
@@
In Figure 13 we show that the qualitative conclusions with regard to the respective procedures are as in § 5.2. Interestingly, in the hardest setting at the bottom row we see that the power of InfoSCOP is even larger than classic conformal. This is because the initial selection step not only prevents paying too much for selection, but it can also improve the calibration set for computing the conformal $p$-values after selection. In this hard setting, there is a large improvement of the tail probability that matters for inference. We demonstrate this in one realization of the data generation in Figure 14, where we can see that the 0.95 quantile of the original calibration set is much larger than the 0.95 quantile of examples that remain in the calibration set after the initial selection step. This is because the data generation is such that examples in the central cloud tend to have larger nonconformity scores $S_{Y_{i}}(X_{i})$ than the examples that are not in the central cloud, and most of the examples from the central cloud were eliminated from the calibration sample with the initial selection step. So the remaining nonconformity scores after the initial selection step tend to be smaller. Since the nonconformity scores of the examples from the test sample are unchanged, the improved calibration set after the initial selection step results in $p$-values that tend to be smaller when testing the null group, and in particular there are many more $p$-values that are below the 0.05 level.
@@
Figure 13: Selecting informative prediction sets in the iid setting with a large overlap between three classes. FCR (left column), resolution-adjusted power (middle column) and the expected fraction of covering prediction sets (right column) versus SNR for: balanced classes for minimally informative prediction sets (first row) and for prediction sets excluding a null class (second row); unbalanced classes for minimally informative prediction sets (third row) and for prediction sets excluding the largest class (fourth row). Each class is a mixture of a common component and a unique component, see § J.2 for details. Based on 2000 data generations, 1000 data points were used for training, and $n=m=500$.
@@
J.3 Three classes of animals
For the class-conditional model with label shift described in § 6, Figure 15 provides the resulting FCR and power measures.
c) Non-trivial classification with label shift
d) Non-null classification with label shift
Figure 15: For the class-conditional model, the FCR, average size of the selected, SR, and resolution-adjusted power for the methods, for $\alpha=0.1$.